OpenStudy (anonymous):

What value of x will make the equation 2x + 6x = 56?

OpenStudy (anonymous):

\[2x+6x=56\]\[8x=56\]\[x=\frac{56}{8}\]\[x=7\]

OpenStudy (anonymous):

2x + 6x = 56 combine like terms 8x = 56 divide both sides by 8 to clear the fraction and to isolate 'x' 8x/8 = 56/8 x = 56/8